5. Decompose


(i)


x2 +x +1


(x +3)(x2 −x +1)



(ii)


x4 −x3 −2x2 +4x +1


x (x −1)2

1
Expert's answer
2022-05-16T16:39:57-0400

(i)


"\dfrac{x^2+x+1}{(x +3)(x^2 −x +1)}=\dfrac{A}{x+3}+\dfrac{Bx+C}{x^2 −x +1}"

"=\dfrac{A(x^2-x+1)+(Bx+C)(x+3)}{(x +3)(x^2 −x +1)}"

"=\dfrac{Ax^2-Ax+A+Bx^2+3Bx+Cx+3C}{(x +3)(x^2 −x +1)}"

"x^2:A+B=1"

"x^1:-A+3B+C=1"

"x^0:A+3C=1"

"A=1-3C"

"B=3C"

"-1+3C+9C+C=1"


"A=\dfrac{7}{13}"

"B=\dfrac{6}{13}"

"C=\dfrac{2}{13}"

"\dfrac{x^2+x+1}{(x +3)(x^2 −x +1)}=\dfrac{\dfrac{7}{13}}{x+3}+\dfrac{\dfrac{6}{13}x+\dfrac{2}{13}}{x^2 −x +1}"


(ii)


"\dfrac{x^4 −x^3 −2x^2 +4x +1}{x(x -1)^2}=\dfrac{x^2(x^2-2x+1)}{x(x -1)^2}"

"+\dfrac{x(x^2-2x+1)}{x(x -1)^2}+\dfrac{-x^2+3x+1}{x(x -1)^2}"

"=x+1+\dfrac{-x^2+3x+1}{x(x -1)^2}"

"\dfrac{-x^2+3x+1}{x(x -1)^2}=\dfrac{A}{x}+\dfrac{B}{x-1}+\dfrac{C}{(x-1)^2}"

"=\dfrac{A(x-1)^2+Bx(x-1)+Cx}{x(x -1)^2}"

"=\dfrac{Ax^2-2Ax+A+Bx^2-Bx+Cx}{x(x -1)^2}"

"x^2:A+B=-1"

"x^1:-2A-B+C=3"

"x^0:A=1"

"A=1"

"B=-2"

"C=3"

"\dfrac{-x^2+3x+1}{x(x -1)^2}=\dfrac{1}{x}+\dfrac{-2}{x-1}+\dfrac{3}{(x-1)^2}"

Therefore

"\dfrac{x^4 −x^3 −2x^2 +4x +1}{x(x -1)^2}"

"=x+1+\dfrac{1}{x}+\dfrac{-2}{x-1}+\dfrac{3}{(x-1)^2}"